# Provable Robustness in Multimodal Large Language Models via Feature Space Smoothing

### v1

@@ 5.2 Experimental results on different tasks
Image captioning. The comparative results of different defense methods and MLLMs on the image captioning task are presented in Table 1. In this task, the adversaries aim to manipulate the MLLMs into producing the caption corresponding to a completely unrelated and maliciously chosen target image. These results demonstrate that converting MLLMs into their smoothed variants via the proposed FS-PSM yields consistently strong robustness across diverse adversarial attacks, while empirical defenses exhibit a noticeable performance drop under stronger attacks. When adversaries vary from M-Attack to FOA, the accuracy of LLaVA with FARE drops from 44% to 19%, and that of TeCoA declines from 51% to 21%. Notably, utilizing FS-PSM achieves remarkable improvements, with the accuracy of LLaVA increasing from 1% to 87% and the ASR dropping from 94% to 1% under the strongest FOA attack. Furthermore, the results demonstrate that FS-PSM possesses strong cross-model generalization; directly integrating it into FARE and TeCoA without any fine-tuning consistently enhances their robustness.
Image classification. The comparative results of different defense methods on the image classification are presented in Table 3. The adversarial objective is to mislead the model into classifying an adversarial image into a maliciously targeted class that is semantically unrelated to the original one. For image classification with CLIP-L14, we train a one-layer classification head on top of its extracted features. The results indicate that FS-PSM provides consistent and substantial robustness improvements for classification tasks. Meanwhile, those adversarial training-based defenses exhibit a noticeable performance drop under strong FOA attacks. The accuracy of FARE with LLaVA drops from 55.4% to 24.8%, and that of TeCoA declines from 67.4% to 29.0% when adversaries vary from M-Attack to FOA. Integrating FS-PSM into the vanilla model achieves remarkable improvements, with the accuracy of LLaVA increasing from 3.8% to 88.0% and the ASR dropping from 81% to 0.4% under the strongest FOA attack. Meanwhile, directly integrating the FS-PSM brings huge robustness gains for those adversarially trained feature encoders.
@@
VQA. The comparative results of different defense methods on LLaVA-1.5-7B on the VQA are presented in Table 4. The performances of OpenFlamingo and CLIP are omitted, as these models are not directly applicable to the VQA task. In this setting, the adversarial objective is to mislead the model into selecting a wrong option, ”None of the above”, for each image–question pair. This task presents a greater challenge for pure vision-based adversaries than the previous two, as MLLMs can often infer correct answers directly from textual cues without relying heavily on visual inputs. The results demonstrate that incorporating our method yields substantial performance improvements across all attack types, significantly enhancing prediction accuracy while reducing the attack success rate to nearly zero. More implementation details for the main results are presented in our supplementary material, section S.3.
5.3 Further analysis on the main results
@@
Reasons for FARE and TeCoA performing poorly with high FCS. Although those encoders preserve high FCS with respect to their clean inputs, the adversarial training process inherently induces a distributional shift in the feature representations. Subsequent adversarial perturbations further amplify this shift, leading to more pronounced performance degradation. In addition, the limited diversity of adversarial training data restricts their feature generalization ability, amplifying the performance drop when evaluated on downstream tasks with unseen or mismatched data distributions.
Advantages of our proposed FS with PSM. The experimental results highlight three key advantages: (1) Effectiveness: utilizing the proposed FS-PSM substantially enhances the performance of various MLLMs under adversarial attacks; (2) Trustworthiness: compared with adversarial training, FS-PSM provides more stable and consistently stronger protection across a wide range of attacks; (3) Generality: FS-PSM can be combined with other defenses to enhance the effectiveness without additional fine-tuning.
5.4 Ablation study
Table 5: The ablation study. We report the average FCSB and the average $\mathcal{R}$ for $\mathrm{FCSB}\geq 0.5$ for certified robustness. We report the average accuracy and ASR for empirical robustness under FOA attack. The clean accuracy without attack is 92.4%.
Encoder $\sigma$ Avg. FCSB at different $\epsilon$ Avg. $\mathcal{R}$ Acc ASR
@@ 5.4 Ablation study
To rigorously assess the contribution of the proposed FS and each module of PSM, we adopt CLIP-B16 as the base feature encoder and evaluate the performance on the image classification task. We report both the theoretically certified robustness and the empirical performance under the strong FOA attack. The results are summarized in Table 5, where the first row indicates the performance of the vanilla CLIP-B16 model without utilizing the FS. The next three rows are the results of the smoothed CLIP-B16 with FS, smoothed CLIP-B16 with FS and mapper $\mathcal{M}$, and smoothed CLIP-B16 with FS and PSM. The results demonstrate that enforcing smoothness in the feature space substantially enhances practical robustness against adversarial perturbations, raising the accuracy from 1.6% to 91.6% and reducing the ASR from 95.6% to 0.4%. Meanwhile, each component makes a significant contribution to improving both the certified radius $\mathcal{R}$ and adversarial robustness.
More experimental results on: 1) the performance of using different purifier (a lightweight U-Net denoiser), 2) the performance under a stronger adversarial bound $\|\epsilon\|_{\infty}=32$, and 3) efficiency–robustness trade-off under various $n_{0}$, are provided in our supplementary material, section S.3.
@@
This work pioneers the research on establishing the certified robustness of MLLMs via a feature-space perspective. By introducing the Feature-space Smoothing (FS) framework, we show how to transform any feature encoder into a smoothed version that is equipped with a theoretical lower bound on the cosine similarity between clean and adversarial representations. Moreover, we introduce the Purifier and Smoothness Mapper (PSM), a plug-and-play module that can be seamlessly integrated into existing MLLMs to amplify the certified robustness under FS without any fine-tuning. Extensive experiments demonstrate that the proposed FS-PSM not only provides strong theoretical robustness guarantees but also achieves superior empirical protection against various attacks.
Supplementary Material
S.7 Proof of core theorem
S.7.1 Proof of lemma1
@@
$\nabla\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))=\frac{\nabla\hat{S}_{x_{c}}(\bm{x})}{\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)}.$ (S.16)
@@
$\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)=\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr)$ (S.17)
@@
$u\cdot\nabla\hat{S}_{{\bm{x}}_{c}}(\bm{x})\leq\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr).$ (S.18)
@@
$\nabla\hat{S}_{x_{c}}(\bm{x})=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}}S_{{\bm{x}}_{c}}(\bm{t})\,({\bm{x}}-{\bm{t}})\,\exp\!\bigl(-\tfrac{1}{2}\|{\bm{x}}-{\bm{t}}\|^{2}\bigr)\,\mathrm{d}{\bm{t}},$ (S.19)
@@
$\mathbb{E}\bigl[(X\cdot u)\,1\{X\cdot u\geq-\Phi^{-1}(p)\}\bigr]=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\tfrac{1}{2}\bigl(\Phi^{-1}(p)\bigr)^{2}\Bigr),$ (S.20)
@@
$\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})).$ (S.21)
@@
$=\tfrac{1}{2}\Big(1+\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle\Big).$ (S.23)
@@
${Cos}\big(\hat{f}({\bm{z}}),\,f_{e}({\bm{x}})\big)$ $=\frac{\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle}{\|\hat{f}({\bm{z}})\|}\;$ (S.25)
@@
$\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})\;\geq\;\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big).$ (S.26)
Combining Equations S.24S.25S.26, we thereby can derive that:
$\mathrm{Cos}\big(\hat{f}({\bm{x}}^{\prime}),\,f_{e}({\bm{x}})\big)\;\geq\;2\,\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big)-1,$ (S.27)
@@
S.8 Training details of PSM
S.8.1 Details of fine-tuning noise purifier
The Purifier $\mathcal{P}$ is optimized in a self-supervised manner using the objective $\mathcal{L}_{p}$ defined in Equation 9. We employ the ImageNet-pretrained $256\times 256$ unconditional diffusion modelS.8.1 as the backbone of $\mathcal{P}$. All input images are resized to $256\times 256$.
To perform denoising, we instantiate a diffusion process consisting of 1,000 timesteps with a linear noise scheduler. Following common practices, the rescaled learned noise variance (i.e., ”rescale_learned_sigma”) is disabled to prevent instability during fine-tuning. We fine-tune the model for a total of 12,000 optimization steps using the AdamW optimizer, with an initial learning rate of $5\times 10^{-5}$. An exponential moving average (EMA) with a decay rate of 0.99 is maintained throughout training to stabilize updates and improve the effective generalization of the learned purifier.
footnotetext: https://github.com/openai/guided-diffusion?tab=readme-ov-file
Figure S.3: Structure of the residual smoothness mapper.
S.8.2 Details of training residual mapper
The structure of the proposed residual smoothness mapper $\mathcal{M}$ is illustrated in Figure S.3. Each block is designed to refine the purified feature representation while preserving its statistical structure and ensuring that the refinement magnitude adapts smoothly to the injected noise magnitude $\sigma$. A single block consists of: (1) noise-aware LayerNorm-FiLM($\sigma$) module, (2) multi-head attention, (3) a channel-wise MLP branch, (4) a depthwise convolution, and (5) a noise-aware residual module.
@@
$\tilde{{\bm{z}}}_{i}=\mathrm{LN}(\tilde{{\bm{z}}}_{i}),\hskip 28.80008pt[\gamma(\sigma),\beta(\sigma)]=\mathrm{FiLM}(\sigma),$ (S.28)
@@
$\mathrm{FiLM}(\sigma)=W_{2}\,\phi(W_{1}\sigma)\in\mathbb{R}^{2D},$ (S.29)
@@
$h_{i}=h_{\text{attn}}+h_{\text{mlp}}+0.5\,h_{\text{conv}},$ (S.30)
@@
$\sigma^{\beta_{i}}=\sigma^{\mathrm{softplus}(\mathrm{MLP}_{\beta}(\sigma))}.$ (S.31)
@@
$\Delta{\bm{z}}_{i}=\left[(1+\gamma(\sigma))\,\bigl(v_{i}\cdot\alpha_{i}\sigma^{\beta_{i}}\bigr)+\beta(\sigma)\right]\odot\mathrm{Scale}_{i},$ (S.32)
@@
$\tilde{{\bm{z}}}_{i+1}=\tilde{{\bm{z}}}_{i}+\Delta{\bm{z}}_{i}.$ (S.33)
@@
Table S.6: Experimental results on adversarial robustness of different defense methods and MLLMs on image classification tasks. The Mapper $\mathcal{M}$ is a lightweight U-Net. The overall best results are shown in bold, and the best results without smoothing are underlined.
Model Method M-Attack [23] FOA [14] AttackVLM [51]
[…]
LLaVA- 1.5-7B Org. 0.427 8.2% 78.2% 0.437 3.8 % 81.0 % 0.458 6.0% 78.4%
Smoothed org. 0.590 84.8% 0.2% 0.600 87.2% 0.2% 0.605 87.0% 0.2%
FARE [37] 0.574 55.4% 14.6% 0.521 24.8% 39.8% 0.508 25.8% 45.0%
Smoothed FARE 0.695 75.6% 5.2% 0.681 66.2% 1% 0.671 68.0% 1%
TeCoA [29] 0.731 67.4% 2% 0.626 29.0% 20.0% 0.592 33.6% 17.8%
Smoothed TeCoA 0.790 70.6% 0.6% 0.714 58.8% 0.8% 0.712 56.4% 0.4%
CLIP -L14 Org. 0.427 18.0% 37.2% 0.437 10.0 % 52.4 % 0.458 16.0% 48.4%
Smoothed org. 0.590 88.0% 0.4% 0.600 90.0% 0.4% 0.604 90.4% 0.4%
FARE [37] 0.574 70.4% 5.2% 0.521 42.0% 14.4% 0.508 35.6% 24.4%
Smoothed FARE 0.695 84.4% 0% 0.682 78.4% 0% 0.671 76.4% 0%
TeCoA [29] 0.731 81.2% 2.4% 0.626 39.2% 20.2% 0.592 35.8% 24.4%
Smoothed TeCoA 0.790 87.6% 1.2% 0.714 75.6% 2.4% 0.712 73.2% 2.0%
Figure S.4: Illustration of the evaluation process.
@@
S.9 More experimental results
S.9.1 Implementation details for main results

### v2

@@ 6.2 Experimental Results on Different Tasks
Image captioning. Table 3 reports the image captioning results of MLLMs under different attacks and defenses. The adversaries aim to manipulate the MLLMs into producing the caption corresponding to a maliciously chosen target image. The results show that converting MLLMs into smoothed variants via the proposed FS smoothing yields consistently strong robustness across diverse attacks, whereas empirical defenses degrade substantially under stronger adversaries. When attacks escalate from M-Attack to FOA, the accuracy of LLaVA with FARE drops from 44% to 19%, and that of TeCoA from 51% to 21%. In contrast, FS achieves significant robustness gains, increasing LLaVA accuracy from 1% to 65% and reducing the ASR from 94% to 6% under the strongest FOA attack. Moreover, FS exhibits strong cross-model generalization, consistently improving the robustness of FARE and TeCoA without additional fine-tuning.
Image classification. Table 4 presents the classification results, indicating that FS delivers consistent and substantial robustness improvements across all settings. The adversarial objective is to mislead the model into classifying an adversarial image into a maliciously targeted class. integrating FS into the vanilla model increases LLaVA accuracy from 3.8% to 87.2% and reduces the ASR from 81% to 0.2% under the strongest FOA attack. Similar robustness gains are observed when FS is applied to FARE and TecoA.
@@
VQA. The comparative results of different defense methods on LLaVA-1.5-7B on the VQA are presented in Table 5. In this setting, the adversarial objective is to mislead the model into selecting a wrong option, ”None of the above”, for each image–question pair. This task presents a greater challenge for pure vision-based adversaries than the previous two, as MLLMs can often infer correct answers directly from textual cues. The results demonstrate that incorporating FS consistently yields substantial performance improvements across all attack types, significantly improving prediction accuracy while driving the ASR to nearly zero.
@@
Reasons for the poor performance of FARE and TeCoA despite high FCS. Although these encoders preserve high FCS relative to their clean inputs, adversarial training inherently induces a shift in the feature-space distribution. Subsequent adversarial perturbations and the limited diversity of adversarial training data further exacerbate this mismatch, leading to pronounced performance degradation on downstream tasks with unseen data distributions.
6.3 Ablation Study and Analysis of Efficiency
Table 6: The ablation study. The clean accuracy is 91.0%.
Smoothing Encoder $\sigma$ Avg. FCSB at different $\epsilon$ Avg. $\mathcal{R}$ Acc
@@ 6.3 Ablation Study and Analysis of Efficiency
[…]
Ablation: To rigorously assess the contribution of the proposed FS and each module of GSB (including denoiser $\mathcal{P}$ and mapper $\mathcal{M}$), we adopt CLIP-L14 as the base feature encoder and conduct the image classification task following the design in Section 5.2. We report both the feature-wise certification and the empirical performance under the FOA attack. The results are summarized in Table 6, where the first row indicates the performance of the vanilla CLIP-L14 without FS. The next four rows are the results of the smoothed CLIP-L14 with different modules enhanced. The results demonstrate that each component makes a significant contribution to improving both the certified radius $\mathcal{R}$ and adversarial robustness. Meanwhile, it demonstrates that enforcing smoothness in the feature space substantially enhances practical robustness against adversarial perturbations, raising the accuracy from 16.0% to 89.6%.
Table 7: The analysis of efficiency.
Model $n_{0}$ Avg. infer time (s) FCS
LLaVA-1.5 7B 1 0.41 0.450
4 0.52 0.562
8 0.69 0.590
64 3.08 0.609
Efficiency analysis: Table 7 presents the efficiency analysis under varying numbers of Gaussian samples, evaluated on a single RTX 4090 GPU. We consider the image captioning task and report the average per-image inference time together with the average FCS under FOA attack. Compared to whole-model smoothing, feature-wise smoothing incurs lower inference overhead; notably, setting $n_{0}=8$ increases the per-image inference time of LLaVA by only $\sim$0.28 seconds, highlighting the practicality of FS for real-world deployment.
@@
This work pioneers the research on establishing the feature-space certified robustness of MLLMs. By introducing the Feature-space Smoothing (FS) framework, we show how to transform a given encoder into a smoothed version that is equipped with a theoretical lower bound on the cosine similarity between clean and adversarial representations. Moreover, we propose the Gaussian Robustness Booster (GSB), a plug-and-play module that seamlessly integrates into existing MLLMs to enhance the certified robustness guaranteed by FS. Extensive experiments demonstrate that FS provides effective adversarial protection for MLLMs in a plug-and-play manner while incurring modest inference latency, making it promising for real-world deployment on MLLM protection.
Appendix
Appendix A.1 Proof of core theorem
A.1.1 Proof of lemma1
@@
$\nabla\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))=\frac{\nabla\hat{S}_{x_{c}}(\bm{x})}{\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)}.$ (A.16)
@@
$\Phi^{\prime}\bigr(\Phi^{-1}(\hat{S}_{x_{c}}(\bm{x}))\bigr)=\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr)$ (A.17)
@@
$u\cdot\nabla\hat{S}_{{\bm{x}}_{c}}(\bm{x})\leq\frac{1}{\sqrt{2\pi}}\exp\!\bigl(-\tfrac{1}{2}(\Phi^{-1}(p))^{2}\bigr).$ (A.18)
@@
$\nabla\hat{S}_{x_{c}}(\bm{x})=\frac{1}{(2\pi)^{n/2}}\int_{\mathbb{R}^{n}}S_{{\bm{x}}_{c}}(\bm{t})\,({\bm{x}}-{\bm{t}})\,\exp\!\bigl(-\tfrac{1}{2}\|{\bm{x}}-{\bm{t}}\|^{2}\bigr)\,\mathrm{d}{\bm{t}},$ (A.19)
@@
$\mathbb{E}\bigl[(X\cdot u)\,1\{X\cdot u\geq-\Phi^{-1}(p)\}\bigr]=\frac{1}{\sqrt{2\pi}}\exp\!\Bigl(-\tfrac{1}{2}\bigl(\Phi^{-1}(p)\bigr)^{2}\Bigr),$ (A.20)
@@ A.1.2 Proof of Theorem1
$\left\|{\bm{x}}-{\bm{x}}^{\prime}\right\|\geq\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}))-\Phi^{-1}(\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})).$ (A.21)
@@
$=\tfrac{1}{2}\Big(1+\langle\hat{f}_{e}({\bm{x}}),\,f_{e}({\bm{x}})\rangle\Big).$ (A.23)
@@
${Cos}\big(\hat{f}({\bm{z}}),\,f_{e}({\bm{x}})\big)$ $=\frac{\langle\hat{f}({\bm{z}}),f_{e}({\bm{x}})\rangle}{\|\hat{f}({\bm{z}})\|}\;$ (A.25)
@@
$\hat{S}_{{\bm{x}}}({\bm{x}}^{\prime})\;\geq\;\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big).$ (A.26)
Combining Equations A.24A.25A.26, we thereby can derive that:
$\mathrm{Cos}\big(\hat{f}({\bm{x}}^{\prime}),\,f_{e}({\bm{x}})\big)\;\geq\;2\,\Phi\!\Big(\Phi^{-1}\!\big(\hat{S}({\bm{x}})\big)-\epsilon\Big)-1,$ (A.27)
@@
Appendix A.2 Training details of GSB
A.2.1 Details of Denoiser
The denoiser $\mathcal{P}$ is implemented as a lightweight conditional U-Net and trained in a supervised denoising paradigm using a mean squared error (MSE) objective to regress clean images from their noisy counterparts. $\mathcal{P}$ adopts a symmetric encoder–decoder architecture composed of residual blocks and multi-scale skip connections, which facilitate effective information propagation across different spatial resolutions. All input images are resized to $256\times 256$, and the network progressively aggregates local and global contextual features through hierarchical downsampling and upsampling operations. To enable adaptive denoising under varying noise levels, $\mathcal{P}$ incorporates a learnable sigma embedding that conditions intermediate feature transformations, allowing the model to dynamically adjust its denoising behavior.
We optimize $\mathcal{P}$ using the Adam optimizer with a learning rate of $1\times 10^{-4}$, a batch size of 8, and a total of 8 training epochs. Notably, $\mathcal{P}$ remains computationally efficient, containing only approximately 10 MB of parameters.
A.2.2 Details of training residual mapper
Figure A.3: Structure of the residual smoothness mapper.
The structure of the proposed residual smoothness mapper $\mathcal{M}$ is illustrated in Figure A.3. Each block is designed to refine the purified feature representation while preserving its statistical structure and ensuring that the refinement magnitude adapts smoothly to the injected noise magnitude $\sigma$. A single block consists of: (1) noise-aware LayerNorm-FiLM($\sigma$) module, (2) multi-head attention, (3) a channel-wise MLP branch, (4) a depthwise convolution, and (5) a noise-aware residual module.
@@ A.2.2 Details of training residual mapper
$\tilde{{\bm{z}}}_{i}=\mathrm{LN}(\tilde{{\bm{z}}}_{i}),\qquad[\gamma(\sigma),\beta(\sigma)]=\mathrm{FiLM}(\sigma),$ (A.28)
@@
$\mathrm{FiLM}(\sigma)=W_{2}\,\phi(W_{1}\sigma)\in\mathbb{R}^{2D},$ (A.29)
@@
$h_{i}=h_{\text{attn}}+h_{\text{mlp}}+0.5\,h_{\text{conv}},$ (A.30)
@@
$\sigma^{\beta_{i}}=\sigma^{\mathrm{softplus}(\mathrm{MLP}_{\beta}(\sigma))}.$ (A.31)
@@
$\Delta{\bm{z}}_{i}=\left[(1+\gamma(\sigma))\,\bigl(v_{i}\cdot\alpha_{i}\sigma^{\beta_{i}}\bigr)+\beta(\sigma)\right]\odot\mathrm{Scale}_{i},$ (A.32)
@@
$\tilde{{\bm{z}}}_{i+1}=\tilde{{\bm{z}}}_{i}+\Delta{\bm{z}}_{i}.$ (A.33)
@@
Appendix A.3 More Details on Prediction-Wise Certification
A.3.1 Preliminaries and Notation
Let $f_{e}:\mathcal{X}\to\mathbb{R}^{d}$ be a (deterministic) feature encoder and let $\hat{f}_{e}$ denote its smoothed version produced by Feature-space Smoothing (FS). Given a clean input ${\bm{x}}\in\mathcal{X}$ and an adversarial input ${\bm{x}}^{\prime}\in\mathcal{X}$ satisfying $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, define the clean feature and the adversarial (smoothed) feature as
${\bm{z}}:=f_{e}({\bm{x}})\in\mathbb{R}^{d},\qquad\hat{\bm{z}}^{\prime}:=\hat{f}_{e}({\bm{x}}^{\prime})\in\mathbb{R}^{d}.$ (A.34)
We further define their $\ell_{2}$-normalized counterparts:
${\bm{u}}:=\frac{{\bm{z}}}{\|{\bm{z}}\|_{2}}\in\mathbb{S}^{d-1},\qquad{\bm{v}}:=\frac{\hat{\bm{z}}^{\prime}}{\|\hat{\bm{z}}^{\prime}\|_{2}}\in\mathbb{S}^{d-1},$ (A.35)
where $\mathbb{S}^{d-1}:=\{{\bm{t}}\in\mathbb{R}^{d}:\|{\bm{t}}\|_{2}=1\}$ is the unit sphere. Thus, $\mathrm{Cos}(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}}))={\bm{u}}^{\top}{\bm{v}}$.
FS-certified feature cosine similarity.
From Theorem 3.2, for any $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, the cosine similarity between $\hat{f}_{e}({\bm{x}}^{\prime})$ and $f_{e}({\bm{x}})$ is lower bounded as
${\bm{u}}^{\top}{\bm{v}}\;=\;\mathrm{Cos}\!\big(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}})\big)\;\geq\;\gamma({\bm{x}};\epsilon),$ (A.36)
where we denote the Feature Cosine Similarity Bound (FCSB) by
$\gamma({\bm{x}};\epsilon)\;:=\;2\Phi\!\left(\Phi^{-1}(\hat{S}({\bm{x}}))-\epsilon\right)-1\in[-1,1].$ (A.37)
For brevity, we write $\gamma:=\gamma({\bm{x}};\epsilon)$.
Cosine-prototype classification head.
We consider a cosine-based prototype classifier on the normalized feature ${\bm{t}}\in\mathbb{S}^{d-1}$. Let $\{{\bm{p}}_{k}\}_{k=1}^{K}\subset\mathbb{R}^{d}$ be learnable class prototypes satisfying $\|{\bm{p}}_{k}\|_{2}=1$ for all $k$. The class score (logit) for class $k$ is defined as
$s_{k}({\bm{u}}):={\bm{u}}^{\top}{\bm{p}}_{k},$ (A.38)
and the predicted label is
$\hat{y}({\bm{u}}):=\arg\max_{k\in\{1,\dots,K\}}s_{k}({\bm{u}})\;=\;\arg\max_{k}{\bm{u}}^{\top}{\bm{p}}_{k}.$ (A.39)
Let $y:=\hat{y}({\bm{u}})$ denote the clean prediction under $\bm{u}$}.
A.3.2 Tight Cosine Similarity Bound
Lemma A.3.1 (Tight bounds under a cosine constraint).
Let ${\bm{u}},{\bm{v}},{\bm{p}}\in\mathbb{S}^{d-1}$ be unit vectors and suppose that ${\bm{u}}^{\top}{\bm{v}}\geq\gamma$ for some $\gamma\in[-1,1]$. Define $a:={\bm{u}}^{\top}{\bm{p}}\in[-1,1]$. Then the inner product ${\bm{v}}^{\top}{\bm{p}}$ is bounded as
$\gamma a-\sqrt{1-\gamma^{2}}\sqrt{1-a^{2}}\;\leq\;{\bm{v}}^{\top}{\bm{p}}\;\leq\;\gamma a+\sqrt{1-\gamma^{2}}\sqrt{1-a^{2}}.$ (A.40)
Proof sketch.
The stated bounds are a standard consequence of spherical geometry. Since ${\bm{u}},{\bm{v}},{\bm{p}}\in\mathbb{S}^{d-1}$, we may interpret their inner products as cosines of the corresponding pairwise angles. In particular, let
$a={\bm{u}}^{\top}{\bm{p}}=\cos\alpha,\qquad\gamma={\bm{u}}^{\top}{\bm{v}}=\cos\Delta,$
where $\alpha,\Delta\in[0,\pi]$ denote the angles between $({\bm{u}},{\bm{p}})$ and $({\bm{u}},{\bm{v}})$, respectively.
Under the constraint ${\bm{u}}^{\top}{\bm{v}}\geq\gamma$, the adversarial feature direction $\bm{v}$} is restricted to lie within a spherical cap centered at $\bm{u}$} with angular radius $\Delta$. Consequently, the angle between $\bm{v}$} and $\bm{p}$} lies in the interval $\alpha-\Delta,\;\alpha+\Delta]$. It follows that
${\bm{v}}^{\top}{\bm{p}}=\cos(\alpha\pm\Delta)=\cos\alpha\cos\Delta\mp\sin\alpha\sin\Delta,$
which yields the bounds in equation A.40 by substituting $\cos\alpha=a$ and $\cos\Delta=\gamma$.
The bounds are tight and are attained when the three vectors ${\bm{u}},{\bm{v}},{\bm{p}}$ lie in a common two-dimensional subspace and $\bm{v}$} is rotated toward or away from $\bm{p}$} along the geodesic on the unit sphere. ∎
A.3.3 Closed-form Propagation to Prediction-level Certification
We now use Lemma A.3.1 to certify that the predicted label remains unchanged under any adversarial perturbation within radius $\epsilon$.
For the clean normalized feature $\bm{u}$}, define the clean class scores
$a_{k}:=s_{k}({\bm{u}})={\bm{u}}^{\top}{\bm{p}}_{k},\qquad k=1,\dots,K.$ (A.41)
Let $y:=\arg\max_{k}a_{k}$ be the clean predicted class.
Theorem A.3.2 (Prediction-level certification via FS with cosine prototypes).
Fix a clean input $\bm{x}$} and let $\gamma=\gamma({\bm{x}};\epsilon)$ be the FS-derived FCSB from Theorem 3.2. For any adversarial input $\bm{x}$$^{\prime}$ satisfying $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$ and each class $k$, we have:
$s_{k}({\bm{v}})={\bm{v}}^{\top}{\bm{p}}_{k}$ $\leq\gamma a_{k}+\sqrt{1-\gamma^{2}}\sqrt{1-a_{k}^{2}},$ (A.42)
$s_{k}({\bm{v}})={\bm{v}}^{\top}{\bm{p}}_{k}$ $\geq\gamma a_{k}-\sqrt{1-\gamma^{2}}\sqrt{1-a_{k}^{2}}.$ (A.43)
In particular, if the following certified margin condition holds:
$\underbrace{\gamma a_{y}-\sqrt{1-\gamma^{2}}\sqrt{1-a_{y}^{2}}}_{=:~\mathrm{LB}_{y}(\gamma)}\;>\;\max_{k\neq y}\underbrace{\Big(\gamma a_{k}+\sqrt{1-\gamma^{2}}\sqrt{1-a_{k}^{2}}\Big)}_{=:~\mathrm{UB}_{k}(\gamma)},$ (A.44)
then the classifier prediction is provably invariant within the $\epsilon$-ball:
$\arg\max_{k}s_{k}({\bm{v}})=y\quad\text{for all }{\bm{x}}^{\prime}\text{ such that }\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon.$ (A.45)
Proof.
By Theorem 3.2, for any $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, we have
$\mathrm{Cos}\!\big(\hat{f}_{e}({\bm{x}}^{\prime}),f_{e}({\bm{x}})\big)={\bm{u}}^{\top}{\bm{v}}\geq\gamma.$ (A.46)
Fix any class $k$. Since ${\bm{u}},{\bm{v}},{\bm{p}}_{k}\in\mathbb{S}^{d-1}$ and ${\bm{u}}^{\top}{\bm{v}}\geq\gamma$, applying Lemma A.3.1 with ${\bm{p}}={\bm{p}}_{k}$ and $a=a_{k}={\bm{u}}^{\top}{\bm{p}}_{k}$ yields the per-class upper and lower bounds equation A.42–equation A.43.
Now consider the clean predicted class $y=\arg\max_{k}a_{k}$. From equation A.43, the adversarial score of class $y$ is bounded from below as
$s_{y}({\bm{v}})\geq\mathrm{LB}_{y}(\gamma).$ (A.47)
For any other class $k\neq y$, from equation A.42 we have
$s_{k}({\bm{v}})\leq\mathrm{UB}_{k}(\gamma).$ (A.48)
If the certified margin condition equation A.44 holds, then for all $k\neq y$,
$s_{y}({\bm{v}})\;\geq\;\mathrm{LB}_{y}(\gamma)\;>\;\mathrm{UB}_{k}(\gamma)\;\geq\;s_{k}({\bm{v}}),$ (A.49)
which implies $s_{y}({\bm{v}})>s_{k}({\bm{v}})$ for every $k\neq y$. Therefore, the argmax prediction under $\bm{v}$} remains $y$, i.e.,
$\arg\max_{k}s_{k}({\bm{v}})=y.$ (A.50)
This holds for any $\bm{x}$$^{\prime}$ with $\|{\bm{x}}^{\prime}-{\bm{x}}\|\leq\epsilon$, which concludes the proof. ∎
Appendix A.4 More Experimental Results
Figure A.4: Illustration of the evaluation process.
A.4.1 Implementation details for main results